OpenStudy (anonymous):

4x-9y=28 3x+25=y solve by substitution

OpenStudy (anonymous):

substitute it in. 4x-9(3x+25)=28

OpenStudy (anonymous):

4x-9(3x+25)=28 -23x-225=28